OpenStudy (anonymous):

Solve the systems using substitution. 6x + 2y = 0 4x + y = -1

OpenStudy (anonymous):

solve the bottom equation for y y=-1-4x

OpenStudy (anonymous):

sub into first eqaution

OpenStudy (anonymous):

6x+2(-1-4x)=0 6x+-2+-8x=0 -2x-2=0 -2x=2 x=-1 solve second equation for when x=-1 4(-1)+y=-1 -4+y=-1 y=3 so your solution set is (x,y)=(-1,3)